### v3

# Sustainable Greenhouse Microclimate Modeling: A Comparative Analysis of Recurrent and Graph Neural Networks

Abstract
The integration of photovoltaic (PV) systems into greenhouses not only optimizes land use but also enhances sustainable agricultural practices by enabling dual benefits of food production and renewable energy generation. However, accurate prediction of internal environmental conditions is crucial to ensure optimal crop growth while maximizing energy production. This study introduces a novel application of Spatio-Temporal Graph Neural Networks (STGNNs) to greenhouse microclimate modeling, comparing their performance with traditional Recurrent Neural Networks (RNNs). While RNNs excel at temporal pattern recognition, they cannot explicitly model the directional relationships between environmental variables. Our STGNN approach addresses this limitation by representing these relationships as directed graphs, enabling the model to capture both environmental dependencies and their directionality. Using high-frequency data collected at 15-minute intervals from a greenhouse in Volos, Greece, we demonstrate that RNNs achieve exceptional accuracy in winter conditions ($R^{2}=0.985$) but show limitations during summer cooling system operation. Though STGNNs currently show lower performance (winter $R^{2}=0.947$), their architecture offers greater potential for integrating additional variables such as PV generation and crop growth indicators.
keywords:
Recurrent neural networks , Spatio-temporal graph neural networks , Graph attention networks , Long short-term memory networks , Greenhouse modeling.
1 Introduction
Modern agriculture faces increasing pressure to optimize land use while reducing its environmental impact. The integration of photovoltaic (PV) systems into agricultural greenhouses has emerged as a promising solution to these challenges, offering dual benefits of food production and renewable energy generation. However, this integration introduces new complexities in greenhouse management, as PV panels directly influence light distribution, temperature patterns, and overall microclimate dynamics. These Agri-PV systems optimize land use by combining solar energy harvesting with protected crop cultivation, offering potential benefits for both energy and food security. Yet the successful implementation of PV-integrated greenhouses requires accurate prediction and control of the internal microclimate to ensure optimal growing conditions while maximizing energy production. While sophisticated physical models incorporating greenhouse thermodynamics, crop physiology, and PV system performance exist, their complexity often restricts real-time applications and rapid design optimization. This study, conducted within the framework of the European REGACE project, represents an initial step toward developing a comprehensive digital twin for PV-integrated greenhouses. Rather than relying on complex physical models, we explore empirical approaches that employ advanced machine learning to predict greenhouse microclimates. The focus on neural network-based models aims to lay the foundation for a more complete digital twin that will ultimately integrate crop growth dynamics and PV energy production. By comparing traditional Recurrent Neural Networks (RNNs) with directed Spatio-Temporal Graph Neural Networks (STGNNs), we investigate whether the explicit modeling of environmental relationships and directional influences can improve prediction accuracy while maintaining computational efficiency. Effective control and optimization of PV-integrated greenhouse environments present a fundamental challenge. Accurate climate control software integrating real-time data from environmental control computers (ECCs) can significantly enhance greenhouse management (Aaslyng et al., 2005). Nonetheless, internal parameters such as temperature and humidity are influenced by various external conditions, including ambient temperature, solar radiation, and wind speed. Conventional modeling approaches can fail to capture the nonlinear and dynamic nature of these relationships, especially when greenhouse data are collected at high frequency (Aaslyng et al., 2003). Researchers have therefore pursued increasingly sophisticated methods to balance these complexities, from physical simulations to data-driven models. Greenhouses represent intricate thermodynamic systems where external conditions and the internal microclimate interplay to influence crop growth and energy use. The introduction of PV panels adds a layer of complexity, prompting further advancements in modeling techniques (Ouazzani Chahidi et al., 2021; Baglivo et al., 2020; Brækken et al., 2023). These approaches range from detailed physics-based studies focusing on ventilation and evapotranspiration (Stanciu et al., 2016; Abdel-Ghany & Kozai, 2006; Mobtaker et al., 2019; Singh et al., 2018; Fitz-Rodríguez et al., 2010) to data-driven or hybrid solutions capable of simulating both microclimate and crop growth processes (Massa et al., 2011). While such models provide essential insights into greenhouse behavior, they can be computationally intensive, and their parameterization often demands multidisciplinary expertise. To address these constraints, researchers have increasingly explored empirical, data-driven approaches that learn greenhouse behavior directly from observations. Early neural network applications demonstrated promise (Seginer et al., 1994; Bussab et al., 2007), evolving into more sophisticated architectures such as Radial Basis Function Neural Networks (Ferreira et al., 2002) and optimized Multi-layer Perceptrons (D’Emilio et al., 2012). A major leap forward came with Recurrent Neural Networks (RNNs), notably Long Short-Term Memory (LSTM) variants, which excel at modeling time-dependent relationships in greenhouse environments (Fourati & Chtourou, 2007; Gharghory, 2020; Hongkang et al., 2018). In parallel, Nonlinear AutoRegressive models with eXogenous inputs (NARX) have shown effectiveness in capturing complex greenhouse system dynamics (Manonmani et al., 2018; Gao et al., 2023). Despite these advances, conventional neural networks predominantly treat environmental variables independently, limiting their ability to depict spatial or directional interconnections. This gap becomes especially relevant in PV-integrated greenhouses, where factors such as solar radiation, humidity, and temperature form intricate environmental interactions that purely temporal models may overlook. Spatio-Temporal Graph Neural Networks (STGNNs) address this shortfall by explicitly modeling intervariable relationships through graph structures (Yu et al., 2018; Sun et al., 2022). Here, each environmental variable becomes a node in a directed graph, capturing how changes in one parameter can propagate to others. The Graph Attention Network (GAT) framework further refines this approach by dynamically weighting node and edge influences, enabling simultaneous modeling of environmental dependencies and temporal evolution (Veličković et al., 2018). Such techniques align with the concept of digital twins in greenhouse systems, where software-based integrative models can guide real-time climate strategies (Aaslyng et al., 2005, 2003) and nutrient management (Massa et al., 2011).
In this study we apply and compare RNNs and directed STGNNs for greenhouse temperature prediction under multiple seasonal conditions. Using data from a Mediterranean greenhouse in Volos, Greece, we assess each model’s predictive accuracy, interpretability via attention mechanisms, and computational feasibility. By evaluating whether directional modeling improves performance over simpler temporal approaches, our work advances the broader REGACE objective of developing efficient, sustainable agrivoltaic greenhouses that integrate microclimate modeling with PV energy production and crop modeling.
The paper is structured as follows: Section 2 describes the data used in this study, including data collection methods, preprocessing steps, and the handling of missing values. Section 3 outlines the methodologies employed, offering a concise introduction to deep learning and detailing the architectures and implementation of the RNN and directed STGNN models. Section 4 presents the analysis and results, comparing the performance of the RNN and directed STGNN models in predicting greenhouse temperature across different seasons. Finally, Sections 5 and 6 discuss the findings, their implications for greenhouse management, and potential directions for future research.
2 Data
The data used in this study originates from a PV greenhouse located in Volos, Greece, for the year 2020, covering from January 9, 2020, to December 21, 2020. The dataset consists of measurements recorded at 15-minute intervals, resulting in 33,446 data points.
The features considered during the analysis are given below. The labels used during the analysis are shown in parentheses:
  • 1.
    External Environmental Factors:
    • (a)
      External Temperature (OUT_temp): The ambient temperature outside the greenhouse.
    • (b)
      External Relative Humidity (OUT_RH): The relative humidity of the air outside the greenhouse.
    • (c)
      Radiation Level (OUT_rad): The amount of solar radiation received.
    • (d)
      Wind Speed (OUT_wind_speed): The speed of the wind outside the greenhouse.
  • 2.
    Internal Environmental Factors:
    • (a)
      Greenhouse Temperature (G2_temp): The internal temperature within Greenhouse.
    • (b)
      Greenhouse Relative Humidity (G2_RH): The internal relative humidity within Greenhouse.
The graphs of the temperature inside and outside the greenhouse for the three periods considered follow:
Figure 1: Temperatures inside and outside the greenhouse for the summer period
Figure 2: Temperatures inside and outside the greenhouse for the autumn period
Figure 3: Temperatures inside and outside the greenhouse for the winter period
To capture the complex relationships between these environmental factors, we represent them as nodes in a directed graph, where edges indicate the direction of influence from one variable to another. The graph is presented in Figure 4.
The directed edges in our graph structure represent fundamental physical relationships between environmental variables, based on well-established thermodynamic and atmospheric physics principles. Our choice of edge directions is motivated by the following physical mechanisms:
  1. 1.
    External Relative Humidity Dependencies:
    • (a)
      Temperature $\rightarrow$ RH: Controls air’s water vapor capacity through the Clausius-Clapeyron relation
    • (b)
      Wind Speed $\rightarrow$ RH: Affects vapor mixing and transport through turbulent diffusion
    • (c)
      Solar Radiation $\rightarrow$ RH: Influences evaporation rates and local temperature gradients
  2. 2.
    External Temperature Dependencies:
    • (a)
      Solar Radiation $\rightarrow$ Temperature: Primary driver through radiative heating
    • (b)
      Wind Speed $\rightarrow$ Temperature: Modifies heat exchange through forced convection
  3. 3.
    Internal Parameters (Temperature and RH) Dependencies:
    • (a)
      External Temperature $\rightarrow$ Internal Temperature: Heat transfer through greenhouse envelope
    • (b)
      External RH $\rightarrow$ Internal RH: Vapor exchange through ventilation and infiltration
    • (c)
      Solar Radiation → Internal Parameters: Direct heating and greenhouse effect
    • (d)
      Wind Speed $\rightarrow$ Internal Parameters: Convective heat exchange rate
    • (e)
      Internal Temperature $\leftrightarrow$ Internal RH: Bidirectional coupling through:
      • i.
        Temperature affecting air’s water-holding capacity
      • ii.
        Humidity influencing evaporative cooling effects
      • iii.
        Plant transpiration processes
These physical relationships form the basis for the graph structure shown in Figure 4, where each edge represents a direct causal influence between variables. This directed graph approach allows our STGNN model to learn and weight these physical dependencies during the prediction process.
Figure 4: Features’ interaction graph
Moreover, the dataset was divided into three seasons to better investigate the seasonal variations in the greenhouse microclimate and due to the high number of missing data. The dates of the divisions are:
  • 1.
    Summer111 During the summer period, the greenhouse employs a cooling system to prevent internal temperatures from rising excessively.: From July 1 to August 31,
  • 2.
    Autumn: From October 9 to November 30,
  • 3.
    Winter: From January 9 to February 29.
We used seasonal names for these periods, even though the dates do not perfectly correspond to the year seasons. This discrepancy is due to a significant number of missing values. For the same reason, spring data were excluded from the analysis due to many missing values recorded between April 23 and April 27.
2.1 Data handling
2.1.1 Missing values
Table 1 shows the number of missing values for each variable in each season.
Table 1: Number of missing values for each season
Season G2_temp G2_RH OUT_temp OUT_RH OUT_rad wind_speed
Summer 0 0 0 0 0 0
Winter 68 68 2 2 2 2
Autumn 75 75 75 75 75 75
To address missing data, we employed a combination of linear interpolation and rolling mean techniques. Linear interpolation is effective for high-density time series data, as it estimates missing values by creating a linear equation between the nearest known data points (Malvar et al., 2004; Taewon et al., 2019), as in Equation 1:
$f(x)=f(x_{o})+\frac{f(x_{1})-f(x_{o})}{x_{1}-x_{o}}(x-x_{o})$ (1)
where $f$ is the linear interpolation function, $x$ the target value and the subscripts 00 and $1$ represent respectively the previous and next value from target value. This approach ensured that the imputed values preserved the continuity and trends in the time series without introducing noticeable distortions or flattening in the data sequences. The rolling mean technique was applied in cases where data exhibited more noise or short-term fluctuations, allowing us to smooth these variations while still accounting for missing values.
2.1.2 Normalization
All features were normalized using the min-max method. This scaling brings all values into the range $[0,1]$, bringing all the variables expressed in different units to the same scale and making them comparable (Alaimo & Seri, 2021).
2.1.3 Train-Test Split
To prepare the data for modeling, each seasonal subset of the dataset was divided into training and testing sets while preserving the chronological order. This approach ensured that the integrity of the time series analysis was maintained, as no future information was inadvertently introduced into the training process. The first 80% of observations within each season were allocated to the training set, while the remaining 20% were used for testing. The resulting date ranges for the training and testing sets in each seasonal period are as follows:
  • 1.
    Winter:
    • (a)
      Training: from 2020-01-09 16:30:00 to 2020-02-19 04:00:00
    • (b)
      Testing: from 2020-02-19 04:15:00 to 2020-02-29 07:00:00
  • 2.
    Summer:
    • (a)
      Training: from 2020-07-01 14:30:00 to 2020-08-19 02:45:00
    • (b)
      Testing: from 2020-08-19 03:00:00 to 2020-08-31 06:00:00
  • 3.
    Autumn:
    • (a)
      Training: from 2020-10-09 22:00:00 to 2020-11-20 13:00:00
    • (b)
      Testing: from 2020-11-20 13:15:00 to 2020-11-30 22:45:00
2.2 Stationarity
In order to make statistical inference based on time series data, it is essential to ensure that the underlying data generative process fulfill the stationary property (Kirchgässner et al., 2012). Stationarity means that the statistical properties of the time series, such as mean and variance, remain consistent over time. This consistency is crucial because it ensures that the patterns observed in the past data are applicable to the future. A time series is stationary if it meets two conditions:
  1. 1.
    The mean (average value) of the series is constant over time.
  2. 2.
    The covariance between two points depends only on the time interval between them, not on their specific positions in time.
Stationarity implies that the series fluctuates around a constant level with consistent variation. This property is essential for making reliable inferences and predictions based on the data (Tsay, 2005). To verify that our data met these criteria, we applied the augmented Dickey-Fuller (ADF) test (Cheung & Lai, 1995). The ADF test results indicate that for all variables in all the three considered periods, the null hypothesis of non-stationarity is rejected, confirming that the series are stationary. The results of the ADF test are presented in Tables 4, 5 and 6 in Appendix A.2.1.
3 Methodology
Neural Networks (NNs) are computational models inspired by the human brain’s neural structure. They consist of interconnected nodes (neurons) organized into layers: an input layer, one or more hidden layers, and an output layer. Each connection between neurons has an associated weight that is adjusted during training to minimize prediction error. Deep Neural Networks (DNNs) extend NNs by incorporating multiple hidden layers, enabling the modeling of complex, nonlinear relationships in data. Activation functions such as ReLU, sigmoid, and tanh introduce nonlinearity, while backpropagation adjusts the weights based on the error rate obtained in previous epochs (Nielsen, 2015).
This section outlines the deep learning techniques employed to develop predictive models of greenhouse temperature. We focus on two architectures: Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs). We discuss their theoretical foundations, how they are applied in this study, and why they are particularly well-suited for modeling greenhouse microclimates.
3.1 Recurrent neural networks
Recurrent Neural Networks (RNNs) are specialized neural networks designed to process sequential data, making them highly suitable for time-series analysis (Jain & Medsker, 1999). Unlike traditional feedforward neural networks, RNNs have recurrent connections that allow information to persist across time steps, enabling the network to capture temporal dependencies.
An RNN processes sequences one element at a time, maintaining a hidden state that captures information about previous inputs. The hidden state $\bm{h}_{t}$ at time step $t$ is updated based on the current input $\bm{x}_{t}$ and the previous hidden state $\bm{h}_{t-1}$:
$\bm{h}_{t}=\tanh(\bm{W}_{h}\bm{x}_{t}+\bm{U}_{h}\bm{h}_{t-1}+\bm{b}_{h})$
The output $\hat{y}_{t}$ at time step $t$ is produced from the hidden state $\bm{h}_{t}$:
$\hat{y}_{t}=\bm{W}_{y}\bm{h}_{t}+\bm{b}_{y}$
where $\bm{W}_{h}$, $\bm{U}_{h}$, and $\bm{W}_{y}$ are weight matrices. $\bm{b}_{h}$ and $\bm{b}_{y}$ are bias vectors. $\tanh$ is the hyperbolic tangent activation function introducing nonlinearity. In deeper RNNs, multiple RNN layers are stacked, passing the hidden state from one layer to the next, allowing the model to capture more complex temporal patterns (Nketiah et al., 2023).
RNNs are effective for modeling how current and past environmental conditions influence future states, making them ideal for predicting variables like greenhouse temperature (Rodriguez et al., 1999). A known challenge with RNNs is the vanishing gradient problem, where gradients diminish during backpropagation through time, making it difficult to learn long-term dependencies (Behrang et al., 2010). However, since our data exhibits short-term dependencies, this issue do not impact our study (Dubinin & Effenberger, 2024).
3.1.1 RNN implementation
We begin by preparing the dataset, where each input sequence $\bm{X}^{(i)}$ consists of historical environmental measurements over a period of $T$ time steps. The features include greenhouse temperature and relative humidity, as well as outdoor environmental variables such as temperature, relative humidity, solar radiation, and wind speed. The target value $y^{(i)}$ is the greenhouse temperature at the next time step.
The notation used in our RNN methodology is defined as follows:
  • 1.
    Datasets and Sequences:
    • (a)
      $\bm{X}^{(i)}\in\mathbb{R}^{T\times F}$: Input sequence for the $i$-th sample, where $T$ is the sequence length and $F$ is the number of features.
    • (b)
      $y^{(i)}\in\mathbb{R}$: Target value for the $i$-th sample.
    • (c)
      $N$: Total number of training samples.
  • 2.
    Model Parameters:
    • (a)
      $E$: Number of training epochs.
    • (b)
      $B$: Batch size.
    • (c)
      $\bm{\theta}$: Set of all learnable parameters in the RNN.
  • 3.
    Layers and Functions:
    • (a)
[…]
    • (b)
      $\text{Dropout}(r)$: Dropout layer with a dropout rate $r$.
    • (c)
      $\text{Dense}(o)$: Fully connected layer with $o$ output units.
    • (d)
      $\text{Loss}(\cdot,\cdot)$: Loss function (Mean Squared Error).
    • (e)
      Optimizer: Optimization algorithm (Adam optimizer).
The RNN model is constructed with two Simple RNN layers, each containing 50 units. The first RNN layer is configured to return sequences (return_sequences=True), allowing the subsequent layer to receive the entire sequence output. Dropout layers with a rate of 0.2 are interleaved between the RNN layers to prevent overfitting by randomly setting a fraction of input units to zero during training. A Dense layer with one unit at the end of the network outputs the predicted temperature.
The model is compiled using the Adam optimizer (Kingma & Ba, 2014), which adapts the learning rate during training for efficient convergence. The loss function is set to Mean Squared Error (MSE), suitable for regression tasks.
Training proceeds for $E$ epochs, where in each epoch, the model is trained on the entire training dataset using batches of size $B$. Optionally, a portion of the training data is set aside for validation to monitor the model’s performance on unseen data during training.
After training, the model is evaluated on the test dataset to assess its predictive performance. Evaluation metrics such as MSE, Root Mean Squared Error (RMSE), and the coefficient of determination ($R^{2}$) are computed to quantify the accuracy of the model’s predictions.
The training algorithm is summarized as follows:
Algorithm 1 Recurrent Neural Network (RNN) Training Algorithm
1:Training sequences $\{(X^{(i)},y^{(i)})\}_{i=1}^{N}$, sequence length $T$, batch size $B$, number of epochs $E$
2:Define model architecture:
3:   Layer 1: Simple RNN with 50 units, input shape $(T,F)$, return sequences True
4:   Layer 2: Dropout layer with rate $0.2$
5:   Layer 3: Simple RNN with 50 units
6:   Layer 4: Dropout layer with rate $0.2$
7:   Layer 5: Dense layer with 1 unit (output)
8:Compile model with optimizer Adam and loss Mean Squared Error
9:for epoch $=1$ to $E$ do
10:     Train the model on training data with batch size $B$
11:     Optionally validate the model on validation data
12:end for
13:Evaluate the model on test data
14:Compute evaluation metrics (MSE, RMSE, $R^{2}$)
3.2 Spatio-Temporal Graph Neural Networks
To model the spatio-temporal dynamics of the greenhouse microclimate, we employ directed STGNN. This approach integrates Graph Attention Networks (GATs) for spatial feature extraction (Veličković et al., 2018) and Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) for temporal sequence modeling. Unlike traditional RNN-based models, which focus primarily on temporal sequences, STGNNs leverage Graph Neural Networks (GNNs) to incorporate topological and relational information among variables. This is well-aligned with recent advances in GNNs for time series (Jin et al., 2024). By representing each environmental variable, such as internal temperature, humidity, and external factors like solar radiation and wind speed, as a node in a directed graph, we capture how changes in one variable may propagate to others through directional edges. Figure 4 illustrates this architecture. Each node corresponds to an environmental variable, while edges encode the direction of influence. At each time step $t$, node features $\bm{X}_{t}$ are extracted from the input sequence $\bm{X}^{(i)}$. To prevent data leakage, the feature of the target node (greenhouse temperature) at the current time step is masked (set to zero) before model input. The GAT layers learn to weigh the importance of edges and neighboring nodes, revealing spatial dependencies, while the LSTM layers capture long-term temporal patterns.
3.2.1 Graph Attention Networks and Directed Graphs
The Graph Attention Network (GAT) layer (Veličković et al., 2018) is employed to handle the directed graph and assign different weights to incoming edges through an attention mechanism. This mechanism allows each node to weigh the importance of its neighbors’ features when updating its own representation.
For each node $i$ and its incoming neighbor $j$, the attention coefficient $\alpha_{ij}$ is computed to determine the importance of node $j$’s features to node $i$:
$e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{W}\bm{x}_{i}\,\|\,\bm{W}\bm{x}_{j}\right]\right)$
$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$
where:
  • 1.
    $\bm{x}_{i}$ and $\bm{x}_{j}$ are the feature vectors of nodes $i$ and $j$, respectively.
  • 2.
    $\bm{W}$ is a learnable weight matrix.
  • 3.
    $\bm{a}$ is a learnable weight vector (attention mechanism).
  • 4.
[…]
  • 5.
    LeakyReLU is the activation function introducing nonlinearity.
  • 6.
    $\mathcal{N}_{i}$ is the set of neighboring nodes sending messages to node $i$.
  • 7.
    $e_{ij}$ is the unnormalized attention score.
  • 8.
    $\alpha_{ij}$ is the normalized attention coefficient, representing the weight assigned to node $j$’s features by node $i$.
The LeakyReLU function, characterized by a small positive slope in its negative part, prevents zero gradients and ensures a more robust learning process, especially important in the attention mechanism where gradient flow is crucial for dynamic weight adjustments.
Node Feature Update:
The updated feature vector $\bm{h}_{i}$ for node $i$ is computed by aggregating the transformed features of its neighbors, weighted by the attention coefficients:
$\bm{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\bm{W}\bm{x}_{j}\right)$
where $\sigma$ is an activation function, such as ReLU.
By using the GAT layer, which inherently supports directed graphs, we effectively model the directionality of interactions among environmental variables. The attention mechanism allows the model to assign different weights to incoming edges based on their importance, capturing the dependence relationships in the data.
3.2.2 Long Short-Term Memory (LSTM) Networks
Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) are a specialized type of RNN architecture designed to better capture and retain long-term dependencies in sequential data. Unlike traditional RNN units, which can suffer from the vanishing gradient problem, LSTM cells incorporate gating mechanisms, namely the input, forget, and output gates, allowing the network to selectively add or remove information from its internal cell state. This structure enables the LSTM to preserve context over many time steps, making it well-suited for complex temporal sequences where relevant patterns may span extended intervals. In the directed STGNN model, the LSTM layer processes the sequence of GAT outputs $\{\bm{H}t\}_{t=1}^{T}$ to extract temporal features that complement the spatial dependencies learned by the GAT layer.
3.2.3 Directed STGNN model implementation
The notation used in our STGNN methodology is defined as follows:
  • 1.
    Datasets and Sequences:
    • (a)
      $\bm{X}^{(i)}\in\mathbb{R}^{T\times N_{n}\times F}$: Input sequence for the $i$-th sample, where $T$ is the sequence length, $N_{n}$ is the number of nodes, and $F$ is the feature size per node (here, $F=1$).
    • (b)
      $y^{(i)}\in\mathbb{R}$: Target value for the $i$-th sample.
    • (c)
      $N$: Total number of training samples.
  • 2.
    Graph Components:
    • (a)
      $E$: Edge index representing directed connections between nodes in the graph.
    • (b)
      $N_{n}$: Number of nodes in the graph.
  • 3.
    Model Parameters:
    • (a)
      $E_{p}$: Number of training epochs.
    • (b)
      $B$: Batch size.
    • (c)
      $H$: Hidden size of the GAT layer.
    • (d)
      $H^{\prime}$: Hidden size of the LSTM layer.
    • (e)
      $K$: Number of attention heads in the GAT layer.
    • (f)
      $\bm{\theta}$: Set of all learnable parameters in the STGNN.
  • 4.
    Layers and Functions:
    • (a)
      $\text{GAT}(F,H,K)$: GAT layer with input feature size $F$, output size $H$ per head, and $K$ attention heads.
    • (b)
      $\text{LSTM}(H\times K\times N_{n},H^{\prime})$: LSTM layer with input size $H\times K\times N_{n}$ and hidden size $H^{\prime}$.
    • (c)
      $\text{Dropout}(r)$: Dropout layer with a dropout rate $r$.
    • (d)
      $\text{Dense}(o)$: Fully connected layer with $o$ output units.
    • (e)
      $\text{Loss}(\cdot,\cdot)$: Loss function (Mean Squared Error).
    • (f)
      Optimizer: Optimization algorithm (Adam optimizer).
Layer 1 (GAT Layer) processes node features $\bm{X}_{t}$ at each time step $t$ using the graph structure defined by $E$. It aggregates information from neighboring nodes based on attention coefficients to capture spatial dependencies. Layer 2 (LSTM Layer) processes the sequence of GAT outputs $\{\bm{H}_{t}\}_{t=1}^{T}$ to capture temporal dependencies over the sequence length $T$. Layer 3 (Dropout Layer) is applied with a rate of 0.2 to mitigate overfitting. Layer 4 (Dense Output Layer) produces the predicted greenhouse temperature $\hat{y}$.
To avoid data leakage, the feature of the target node (greenhouse temperature) at the current time step is set to zero during input preparation. The model is compiled using the Adam optimizer (Kingma & Ba, 2014) and the Mean Squared Error (MSE) loss function. It is trained over $E_{p}$ epochs using batches of size $B$. During training, gradients are computed through backpropagation, and model parameters $\bm{\theta}$ are updated accordingly.
After training, the model is evaluated on the test dataset. Evaluation metrics such as MSE, RMSE, and $R^{2}$ score are calculated to assess the model’s performance.
The training algorithm is summarized as follows:
Algorithm 2 Directed STGNN Training Algorithm
1:Training sequences $\{(\bm{X}^{(i)},y^{(i)})\}_{i=1}^{N}$, edge index $E$, sequence length $T$, batch size $B$, number of epochs $E_{p}$, number of nodes $N_{n}$
2:Initialize model parameters $\bm{\theta}$
3:Define model architecture:
4:   Layer 1: $\text{GAT}(F,H,K)$
5:   Layer 2: $\text{LSTM}(H\times K\times N_{n},H^{\prime})$
6:   Layer 3: $\text{Dropout}(0.2)$
7:   Layer 4: $\text{Dense}(1)$ (output layer)
8:for epoch $=1$ to $E_{p}$ do
9:     for each batch $\{(\bm{X}_{\text{batch}},y_{\text{batch}})\}$ do
10:         Forward pass:
11:         for time step $t=1$ to $T$ do
12:              Extract node features $\bm{X}_{t}$ from $\bm{X}_{\text{batch}}$ at time $t$
13:              Prevent data leakage: Set target node feature $\bm{x}_{v}\leftarrow 0$
14:              Apply GAT layer: $\bm{H}_{t}\leftarrow\text{GAT}(\bm{X}_{t},E)$
15:         end for
16:         Stack $\bm{H}_{t}$ over time to form $\bm{H}_{\text{seq}}$
17:         Apply LSTM layer: $\bm{H}_{\text{lstm}}\leftarrow\text{LSTM}(\bm{H}_{\text{seq}})$
18:         Apply Dropout: $\bm{H}_{\text{drop}}\leftarrow\text{Dropout}(\bm{H}_{\text{lstm}})$
19:         Compute output: $\hat{y}\leftarrow\text{Dense}(\bm{H}_{\text{drop}})$
20:         Compute loss: $L\leftarrow\text{Loss}(y_{\text{batch}},\hat{y})$
21:         Backward pass: Compute gradients $\nabla_{\bm{\theta}}L$
22:         Update parameters: $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}L$
23:     end for
24:     Optionally validate the model on validation data
25:end for
26:Evaluate the model on test data
27:Compute evaluation metrics (MSE, RMSE, $R^{2}$)
4 Analysis and results
This section presents the results obtained from applying the RNN and directed STGNN models to the greenhouse temperature prediction task. Both models were trained and tested using a sequence length of $T=96$ time steps, corresponding to a full 24-hour cycle of observations recorded at 15-minute intervals. This setup allowed the models to capture daily patterns and short-term dependencies effectively.
In addition, both models were trained for 32 epochs with a batch size of 96, aligning with practices in comparable studies (Belhaj Salah & Fourati, 2021; Dae-Hyun et al., 2020). Furthermore, 10% of the training set in each seasonal period was used as a validation set to monitor the models’ performance during training. This approach prevented overfitting, evidenced by the validation loss never surpassing the training loss throughout the process.
To facilitate direct comparisons between the RNN and STGNN models, the same evaluation metrics were employed. Mean Squared Error (MSE) and Root Mean Squared Error (RMSE) measured the prediction error magnitude, while the coefficient $R^{2}$ assessed the proportion of variance in greenhouse temperature explained by the model’s predictions.
4.1 Recurrent neural networks application
Table 2 summarizes the test set results obtained by the RNN model for the three considered seasonal periods.
Table 2: RNN test performance across different seasons
Season MSE RMSE $\bm{R^{2}}$
Summer 0.0032 0.056 0.937
Winter 0.0005 0.022 0.985
Autumn 0.0007 0.026 0.983
The RNN demonstrates strong predictive capabilities, achieving high $R^{2}$ values across all seasons. Its performance is particularly impressive in winter, where $R^{2}\approx 0.985$ indicates that the model captures nearly all the variance in greenhouse temperature. This result depends by several factors related to greenhouse operation during the cold season. First, despite winter typically exhibiting large external temperature fluctuations, the greenhouse’s effectively maintains stable internal conditions, making the system’s behavior more predictable. Second, the reduced solar radiation intensity in winter means the greenhouse is less subject to sudden heat gains that could perturb the internal environment. Third, the thermal screen’s regular nighttime deployment (activated whenever the outside temperature is 2°C lower than the heating setpoint) creates a more controlled environment with reduced thermal exchanges. These factors combine to create a more deterministic system behavior that the RNN can effectively learn and predict, despite the external environment’s inherent variability.
In contrast, while the summer period still exhibits a high $R^{2}$ of about $0.937$, the prediction errors (MSE and RMSE) are relatively higher, with distinct patterns visible in the model’s performance. As shown in Figure 9 the model’s predictions deviate significantly from observations at higher temperatures, producing a characteristic box-shaped pattern in the scatter plot. This behavior directly results from the greenhouse’s cooling system activation, which triggers when internal temperature reaches a threshold of 25°C. Since our empirical model does not explicitly account for this control system, it fails to capture the sudden temperature regulation that occurs above this threshold. However, examining the lower portion of the scatter plot (below normalized temperature values of 0.6), where the cooling system remains inactive, reveals that the model maintains excellent predictive accuracy. This dichotomy in performance clearly demonstrates how the absence of cooling system dynamics in our model impacts its predictive capabilities, particularly during periods of peak thermal stress when temperature control mechanisms are most active.. Autumn performance falls between these two cases, with $R^{2}\approx 0.983$, indicating strong predictive accuracy albeit slightly less than that seen in winter.
Figure 5 provides a scatter plot comparing observed and predicted greenhouse temperatures during the winter period. Points closely aligned along the 1:1 line signify accurate predictions. Figure 6 shows the time series of actual versus predicted values, illustrating how well the RNN’s predictions track the observed temperatures over time.
Figure 5: Scatter plot of the observed vs. predicted greenhouse temperature for the winter period using the RNN model.
Figure 6: Time series of the observed and predicted greenhouse temperature for the winter period using the RNN model.
Similar plots for summer and autumn are included in Appendix A.1.1. These additional figures reveal analogous tendencies: the model accurately tracks observed temperatures, especially in winter, where conditions appear more stable, allowing the RNN to predict internal temperature with remarkable confidence.
While the RNN achieves robust performance in all three periods, the differences across seasons reveal important insights about greenhouse control systems and model limitations. The near-ideal predictions in winter reflect the greenhouse’s effective temperature stabilization, creating a more predictable environment despite external variability. Autumn shows similar high performance, as temperatures rarely reach extremes requiring intensive control interventions. However, summer predictions show a clear dichotomy: excellent accuracy below the cooling threshold temperature, but significant deviations when the cooling system activates at 25°C. This pattern demonstrates that our empirical model’s performance is strongly tied to the greenhouse’s control regime rather than external conditions alone, performing best when passive thermal regulation dominates and struggling when active cooling interventions occur.
4.2 Spatio-Temporal Graph Neural Network application
Table 3 presents the directed STGNN test set results for the three seasonal periods, using the same evaluation metrics as applied to the RNN model. As before, the sequence length, training epochs, batch size, and validation strategy remain consistent.
Table 3: STGNN test performance across different seasons
Season MSE RMSE $\bm{R^{2}}$
Summer 0.0061 0.078 0.887
Winter 0.0023 0.048 0.947
Autumn 0.0034 0.058 0.904
While the STGNN model still achieves relatively high $R^{2}$ values, its overall performance is lower than that of the RNN. In winter, the STGNN attains an $R^{2}$ of approximately 0.947, which, although strong, falls short of the near-perfect predictions obtained by the RNN. Similarly, in summer and autumn, the STGNN yields lower $R^{2}$ scores and higher MSE and RMSE values than the RNN, indicating that incorporating spatial dependencies and directionality did not enhance predictive accuracy in this particular setting.
Figure 7 provides a scatter plot of observed versus predicted temperatures for the winter period, and Figure 8 shows the corresponding time series comparison. These visualizations confirm that, while the STGNN’s predictions track the overall trends, the model does not reach the same level of precision demonstrated by the RNN.
Figure 7: Scatter plot of observed vs. predicted greenhouse temperature for the winter period using the STGNN model.
Figure 8: Time series of observed and predicted greenhouse temperature for the winter period using the STGNN model.
Plots for summer and autumn are included in Appendix A.1.2. They similarly show that the STGNN model, despite considering spatial relationships and directed interactions, does not outperform the simpler temporal approach employed by the RNN. This outcome suggests that, for the given data and conditions, capturing complex spatial-temporal dependencies does not necessarily translate into improved predictive accuracy.
5 Discussion
The comparative evaluation of RNN and directed STGNN models for greenhouse temperature prediction provides insights into how each model leverages available information and handles environmental complexity. The strong performance of the RNN, particularly during the winter period suggests that temporal patterns alone can effectively capture many of the dependencies underlying the greenhouse microclimate. Under these circumstances, long-term stationarity and moderate seasonal variations allow historical sequences of internal and external factors to sufficiently inform future conditions. As a result, the RNN consistently achieves near-perfect predictions in such stable regimes.
However, as conditions grow more complex, the advantages of a temporal-only model appear less pronounced. In summer, for example, the accuracy of the RNN decreases, indicating the challenges posed by more dynamic and rapidly fluctuating environmental factors. Abrupt changes in solar radiation, the intermittent operation of cooling systems, or sudden shifts in external temperature highlight a broader limitation: the RNN’s inability to explicitly model the spatial and directional interactions among variables. While it excels in extracting temporal dependencies, it cannot directly encode the influence structure that might clarify why certain variables cause distinct, localized changes or how perturbations propagate from one factor to another.
In principle, the STGNN architecture is designed to address this very limitation. By representing each variable as a node in a directed graph and capturing the directionality of interactions, the STGNN should be well-suited to handle scenarios where spatial relationships and directional effects matter. The GAT component of the STGNN aims to highlight influential nodes and edges, while the LSTM component extracts temporal dependencies. Together, this framework has the potential to disentangle complex spatio-temporal patterns that a purely temporal model may overlook.
Yet, in our experiments, the STGNN did not outperform the RNN. This outcome can be attributed to several factors. First, the set of input variables, though representative of basic external and internal greenhouse conditions, may not be rich enough to fully leverage the STGNN’s spatial modeling capabilities. For instance, without additional features that intensify spatial heterogeneity (e.g., PV generation data influencing localized temperature gradients, or crop growth indicators that reflect biological processes tied to specific sections of the greenhouse), the STGNN may not find compelling reasons to depart from solutions similar to those of a temporal-only model. In other words, if the underlying system appears nearly homogeneous from a spatial perspective, or if directional edges do not provide sufficiently distinct pathways of influence, the added complexity of the STGNN architecture may not yield tangible benefits.
Second, the STGNN’s performance depends on accurately learning graph structures and attention patterns that differentiate the importance of certain nodes and edges. If the graph connections are not strongly differentiated, the GAT may not produce markedly different node embeddings, diminishing the model’s advantage. The LSTM layer, while capable of capturing temporal information, cannot compensate if the spatial relationships remain underexploited.
These observations guide the interpretation of our results. The theoretical superiority of STGNN in handling complex and directed spatio-temporal dependencies is not dismissed by the current findings. Instead, the experiments suggest that to unlock the STGNN’s full potential, the modeling scenario must present more pronounced spatial variations, clearer directional influences, or a richer set of variables that intensify these spatial-temporal complexities. Under such enriched conditions, the STGNN may provide insights and accuracies that temporal-only models cannot match.
6 Conclusions and future developments
This study examined the performance of Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs) in predicting greenhouse temperature under different seasonal conditions. Despite the STGNN’s conceptual ability to model spatial and directional relationships, the RNN model yielded higher predictive accuracy given the current set of input variables and operating conditions. The strong performance of the RNN, particularly during the winter period, illustrates that purely temporal modeling can effectively capture short-term dependencies.
The implications of this research extend far beyond technical model comparisons, touching on critical aspects of sustainable agriculture and efficient land use. Predicting greenhouse microclimate accurately is a crucial first step toward integrating photovoltaic (PV) systems into greenhouses. This integration promises to optimize land use by combining energy production with food cultivation, thereby enhancing sustainable agricultural practices. It is instrumental in promoting sustainable land management practices that are vital in addressing global environmental challenges, such as resource efficiency and climate change mitigation.
Looking ahead, a more diverse set of input variables may enhance the potential benefits of incorporating spatial and directional information. Despite the STGNN’s current performance not surpassing the RNN in this specific scenario, its architectural advantages position it as a promising framework for future greenhouse modeling challenges. As we move toward integrating photovoltaic systems, the STGNN’s ability to handle directed relationships will become crucial for modeling how PV panels affect local shading patterns, temperature gradients, and energy flows. The graph structure naturally accommodates the addition of new nodes representing PV generation parameters, inverter efficiency, and panel temperature, along with their complex interactions with the greenhouse environment. Similarly, when incorporating crop models, the STGNN can represent different plant varieties as separate nodes, each with unique growth parameters and responses to environmental conditions, enabling simulation of mixed-crop scenarios within the same greenhouse. This extensibility last to potential future additions such as energy storage systems, artificial lighting, or multiple greenhouse compartments with distinct microclimates. Furthermore, the attention mechanism in GAT layers could prove valuable in identifying which relationships become more or less important under different operating conditions or seasonal changes, providing insights that could inform greenhouse design and control strategies. This scalability and interpretability make the STGNN framework particularly valuable for developing comprehensive digital twins of increasingly complex agricultural systems, even if simpler models may suffice for basic temperature prediction tasks.
Such tests could confirm whether additional complexity and diversity in data naturally lead to scenarios where spatial-temporal modeling offers clear advantages, ultimately guiding the development of more robust and universally applicable predictive solutions for controlled-environment agriculture.
This study represents an initial step in the broader REGACE project’s vision of developing comprehensive digital twins for PV-integrated greenhouses. While our current models focus on temperature prediction, they lay the groundwork for REGACE’s more ambitious goals. The project aims to create an integrated modeling framework that combines greenhouse microclimate prediction with PV system performance and crop growth dynamics. Future work within REGACE will focus on expanding these models to incorporate PV panel characteristics, testing different panel configurations and their impact on both energy generation and crop growth. Additionally, the project will explore how these empirical models can complement detailed physical simulations, potentially offering rapid preliminary assessments for greenhouse design optimization. By developing these tools within the REGACE framework, we aim to support the widespread adoption of agrivoltaic greenhouse systems while ensuring optimal conditions for both energy generation and crop production.
7 Acknowledgment
This research was carried out within the framework of the European REGACE project, funded by the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101096056.
Appendix A Supplementary Information
A.1 Figures
A.1.1 RNN results figures
The graphs of the RNN results for the summer and autumn periods follow.
Figure 9: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the RNN model.
Figure 10: Time series of the observed and predicted greenhouse temperature for the summer period using the RNN model.
Figure 11: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the RNN model.
Figure 12: Time series of the observed and predicted greenhouse temperature for the autumn period using the RNN model.
A.1.2 STGNN temperature figures
The graphs of the directed STGNN results for the summer and autumn periods follow.
Figure 13: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the STGNN model.
Figure 14: Time series of the observed and predicted greenhouse temperature for the summer period using the STGNN model.
Figure 15: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the STGNN model.
Figure 16: Time series of the observed and predicted greenhouse temperature for the autumn period using the STGNN model.
A.2 Supplementary tables
A.2.1 ADF test results tables
Hypotheses
  • 1.
    $\mbox{H}_{0}$: The series has a unit root (non-stationary),
  • 2.
    $\mbox{H}_{1}$: The series is stationary.
Table 4: ADF test results for stationarity for the summer period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -15.855 <0.01 Stationary
G2_RH -11.471 <0.01 Stationary
OUT_temp -21.311 <0.01 Stationary
OUT_RH -16.621 <0.01 Stationary
OUT_rad -21.836 <0.01 Stationary
OUT_wind_speed -15.131 <0.01 Stationary
Table 5: ADF test results for stationarity for the autumn period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -16.11 <0.01 Stationary
G2_RH -13.328 <0.01 Stationary
OUT_temp -14.324 <0.01 Stationary
OUT_RH -15.253 <0.01 Stationary
OUT_rad -17.418 <0.01 Stationary
OUT_wind_speed -10.395 <0.01 Stationary
Table 6: ADF test results for stationarity for the winter period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -18.623 <0.01 Stationary
G2_RH -14.903 <0.01 Stationary
OUT_temp -13.869 <0.01 Stationary
OUT_RH -13.673 <0.01 Stationary
OUT_rad -17.356 <0.01 Stationary
OUT_wind_speed -9.815 <0.01 Stationary

### v4

@@
 Our STGNN approach addresses this limitation by representing these relationships as directed graphs, enabling the model to capture both environmental dependencies and their directionality.
-Using high-frequency data collected at 15-minute intervals from a greenhouse in Volos, Greece, we demonstrate that RNNs achieve exceptional accuracy in winter conditions ($R^{2}=0.985$) but show limitations during summer cooling system operation.
-Though STGNNs currently show lower performance (winter $R^{2}=0.947$), their architecture offers greater potential for integrating additional variables such as PV generation and crop growth indicators.
+We benchmark RNNs against directed STGNNs on two 15-min-resolution datasets from Volos (Greece): a six-variable 2020 installation and a more complex eight-variable greenhouse monitored in autumn 2024.
+In the simpler 2020 case the RNN attains near-perfect accuracy, outperforming the STGNN.
+When additional drivers are available in 2024, the STGNN overtakes the RNN ($R^{2}=0.905$ vs $0.740$), demonstrating that explicitly modelling directional dependencies becomes critical as interaction complexity grows.
+These findings indicate when graph-based models are warranted and provide a stepping-stone toward digital twins that jointly optimise crop yield and PV power in agrivoltaic greenhouses.
 keywords:
 Recurrent neural networks , Spatio-temporal graph neural networks , Graph attention networks , Long short-term memory networks , Greenhouse modeling.
+{highlights}
+Compare RNNs and directed STGNNs for greenhouse microclimate prediction.
+RNN excels in the simple 2020 scenario.
+STGNN beats RNN on a complex 2024 greenhouse with eight environmental drivers.
+Directional graphs help when added variables amplify interaction paths.
+Result guides when to pick temporal vs. spatio-temporal models.
 1 Introduction
@@
 Yet the successful implementation of PV-integrated greenhouses requires accurate prediction and control of the internal microclimate to ensure optimal growing conditions while maximizing energy production.
-While sophisticated physical models incorporating greenhouse thermodynamics, crop physiology, and PV system performance exist, their complexity often restricts real-time applications and rapid design optimization.
+While sophisticated physical models incorporating greenhouse thermodynamics, crop physiology, and PV system performance exist [1], their complexity often restricts real-time applications and rapid design optimization.
 This study, conducted within the framework of the European REGACE project, represents an initial step toward developing a comprehensive digital twin for PV-integrated greenhouses.
 Rather than relying on complex physical models, we explore empirical approaches that employ advanced machine learning to predict greenhouse microclimates.
-The focus on neural network-based models aims to lay the foundation for a more complete digital twin that will ultimately integrate crop growth dynamics and PV energy production.
-By comparing traditional Recurrent Neural Networks (RNNs) with directed Spatio-Temporal Graph Neural Networks (STGNNs), we investigate whether the explicit modeling of environmental relationships and directional influences can improve prediction accuracy while maintaining computational efficiency.
-Effective control and optimization of PV-integrated greenhouse environments present a fundamental challenge.
-Accurate climate control software integrating real-time data from environmental control computers (ECCs) can significantly enhance greenhouse management (Aaslyng et al., 2005).
-Nonetheless, internal parameters such as temperature and humidity are influenced by various external conditions, including ambient temperature, solar radiation, and wind speed.
-Conventional modeling approaches can fail to capture the nonlinear and dynamic nature of these relationships, especially when greenhouse data are collected at high frequency (Aaslyng et al., 2003).
-Researchers have therefore pursued increasingly sophisticated methods to balance these complexities, from physical simulations to data-driven models.
-Greenhouses represent intricate thermodynamic systems where external conditions and the internal microclimate interplay to influence crop growth and energy use.
-The introduction of PV panels adds a layer of complexity, prompting further advancements in modeling techniques (Ouazzani Chahidi et al., 2021; Baglivo et al., 2020; Brækken et al., 2023).
-These approaches range from detailed physics-based studies focusing on ventilation and evapotranspiration (Stanciu et al., 2016; Abdel-Ghany & Kozai, 2006; Mobtaker et al., 2019; Singh et al., 2018; Fitz-Rodríguez et al., 2010) to data-driven or hybrid solutions capable of simulating both microclimate and crop growth processes (Massa et al., 2011).
-While such models provide essential insights into greenhouse behavior, they can be computationally intensive, and their parameterization often demands multidisciplinary expertise.
-To address these constraints, researchers have increasingly explored empirical, data-driven approaches that learn greenhouse behavior directly from observations.
-Early neural network applications demonstrated promise (Seginer et al., 1994; Bussab et al., 2007), evolving into more sophisticated architectures such as Radial Basis Function Neural Networks (Ferreira et al., 2002) and optimized Multi-layer Perceptrons (D’Emilio et al., 2012).
-A major leap forward came with Recurrent Neural Networks (RNNs), notably Long Short-Term Memory (LSTM) variants, which excel at modeling time-dependent relationships in greenhouse environments (Fourati & Chtourou, 2007; Gharghory, 2020; Hongkang et al., 2018).
-In parallel, Nonlinear AutoRegressive models with eXogenous inputs (NARX) have shown effectiveness in capturing complex greenhouse system dynamics (Manonmani et al., 2018; Gao et al., 2023).
-Despite these advances, conventional neural networks predominantly treat environmental variables independently, limiting their ability to depict spatial or directional interconnections.
-This gap becomes especially relevant in PV-integrated greenhouses, where factors such as solar radiation, humidity, and temperature form intricate environmental interactions that purely temporal models may overlook.
-Spatio-Temporal Graph Neural Networks (STGNNs) address this shortfall by explicitly modeling intervariable relationships through graph structures (Yu et al., 2018; Sun et al., 2022).
-Here, each environmental variable becomes a node in a directed graph, capturing how changes in one parameter can propagate to others.
-The Graph Attention Network (GAT) framework further refines this approach by dynamically weighting node and edge influences, enabling simultaneous modeling of environmental dependencies and temporal evolution (Veličković et al., 2018).
-Such techniques align with the concept of digital twins in greenhouse systems, where software-based integrative models can guide real-time climate strategies (Aaslyng et al., 2005, 2003) and nutrient management (Massa et al., 2011).
-In this study we apply and compare RNNs and directed STGNNs for greenhouse temperature prediction under multiple seasonal conditions.
-Using data from a Mediterranean greenhouse in Volos, Greece, we assess each model’s predictive accuracy, interpretability via attention mechanisms, and computational feasibility.
-By evaluating whether directional modeling improves performance over simpler temporal approaches, our work advances the broader REGACE objective of developing efficient, sustainable agrivoltaic greenhouses that integrate microclimate modeling with PV energy production and crop modeling.
+By comparing traditional Recurrent Neural Networks (RNNs) with directed Spatio-Temporal Graph Neural Networks (STGNNs), we investigate whether explicit modeling of environmental relationships and directional influences can improve prediction accuracy while maintaining computational efficiency.
+RNNs capture temporal patterns but treat each variable independently, whereas STGNNs embed each environmental variable as a node in a directed graph, allowing the model to learn how changes propagate through explicit causal pathways [2, 3, 4].
+We analyse 15-min data from two Mediterranean greenhouses in Volos (Greece): GH2 (2020, six variables, $\approx$33 k samples across three seasons) and the more complex GH4 (autumn 2024, eight variables including PAR and CO2, 4 k training / 1 k test).
+Both models use identical hyperparameter budgets and a 96-step (24 h) input window, ensuring that performance differences arise from the modelling paradigm rather than tuning effort.
+A vanilla RNN achieves near-perfect accuracy on GH2, yet the STGNN overtakes it on GH4 ($R^{2}=0.905$ vs 0.740), demonstrating that directional graphs earn their keep when additional sensors amplify interaction pathways.
+This finding addresses the open question of when graph machinery is necessary for greenhouse prediction, complementing earlier empirical studies that relied solely on temporal models [5, 6, 7, 8, 9, 10, 11, 12].
+It also advances digital-twin research for PV-integrated houses by showing how physics-informed edge sets can improve data-driven predictions without exhaustive search.
+Effective micro-climate control remains challenging because internal temperature and humidity respond non-linearly to ambient weather, solar radiation and ventilation events [13, 14], and PV panels introduce further heterogeneity [15, 16, 17].
+Hybrid, physics-based simulations capture those processes [18, 19, 20, 21, 22, 23] but are often too slow for on-line use; our results show that a graph-augmented neural network can bridge the gap when variable interactions are rich.
 The paper is structured as follows: Section 2 describes the data used in this study, including data collection methods, preprocessing steps, and the handling of missing values.
@@
 2 Data
-The data used in this study originates from a PV greenhouse located in Volos, Greece, for the year 2020, covering from January 9, 2020, to December 21, 2020.
-The dataset consists of measurements recorded at 15-minute intervals, resulting in 33,446 data points.
+This study draws on two high-frequency (15-min) data sets collected in Volos (Greece) under the REGACE project.
+GH2 spans an entire year (January to December 2020; 33 446 records) with six environmental variables, whereas GH4 is an autumn 2024 campaign (9 October to 20 November 2024; 5 000 records) that adds PAR and CO2 sensors.
+GH2 (2020) – six-variable benchmark
 The features considered during the analysis are given below. The labels used during the analysis are shown in parentheses:
@@
       Greenhouse Relative Humidity (G2_RH): The internal relative humidity within Greenhouse.
-The graphs of the temperature inside and outside the greenhouse for the three periods considered follow:
-Figure 1: Temperatures inside and outside the greenhouse for the summer period
-Figure 2: Temperatures inside and outside the greenhouse for the autumn period
-Figure 3: Temperatures inside and outside the greenhouse for the winter period
-To capture the complex relationships between these environmental factors, we represent them as nodes in a directed graph, where edges indicate the direction of influence from one variable to another.
-The graph is presented in Figure 4.
+The inside-outside temperature trajectories for the autumn period in GH2 is shown in Figure 1.
+The winter and summer periods for GH2 are shown in 9 and 11 in Appendix A.
+During the summer period, the greenhouse employs a cooling system to prevent internal temperatures from rising excessively.
+Spring was excluded because of the high number of missing values.
+For modelling each season is split chronologically: the first 80 % of records form the training set, the remaining 20 % the test set.
+Figure 1: GH2 temperatures inside and outside the greenhouse for the autumn period
 The directed edges in our graph structure represent fundamental physical relationships between environmental variables, based on well-established thermodynamic and atmospheric physics principles.
@@
         Plant transpiration processes
-These physical relationships form the basis for the graph structure shown in Figure 4, where each edge represents a direct causal influence between variables.
+These physical relationships form the basis for the graph structure shown in Figure 2, where each edge represents a direct influence between variables.
 This directed graph approach allows our STGNN model to learn and weight these physical dependencies during the prediction process.
-Figure 4: Features’ interaction graph
-Moreover, the dataset was divided into three seasons to better investigate the seasonal variations in the greenhouse microclimate and due to the high number of missing data.
-The dates of the divisions are:
+Figure 2: GH2 feature-interaction graph
+GH4 (2024) – eight-variable greenhouse
+Compared with GH2 the 2024 installation is instrumented with two additional physical dimensions: photosynthetically active radiation (PAR) and carbon-dioxide concentration, both measured inside and outside the structure:
   • 1.
-    Summer111 During the summer period, the greenhouse employs a cooling system to prevent internal temperatures from rising excessively.: From July 1 to August 31,
+    (G4_PAR) / (OUT_PAR): PAR is the spectral band plants use for photosynthesis.
   • 2.
-    Autumn: From October 9 to November 30,
-  • 3.
-    Winter: From January 9 to February 29.
-We used seasonal names for these periods, even though the dates do not perfectly correspond to the year seasons.
-This discrepancy is due to a significant number of missing values.
-For the same reason, spring data were excluded from the analysis due to many missing values recorded between April 23 and April 27.
-2.1 Data handling
-2.1.1 Missing values
-Table 1 shows the number of missing values for each variable in each season.
-Table 1: Number of missing values for each season
-Season G2_temp G2_RH OUT_temp OUT_RH OUT_rad wind_speed
-Summer 0 0 0 0 0 0
-Winter 68 68 2 2 2 2
-Autumn 75 75 75 75 75 75
-To address missing data, we employed a combination of linear interpolation and rolling mean techniques.
-Linear interpolation is effective for high-density time series data, as it estimates missing values by creating a linear equation between the nearest known data points (Malvar et al., 2004; Taewon et al., 2019), as in Equation 1:
-$f(x)=f(x_{o})+\frac{f(x_{1})-f(x_{o})}{x_{1}-x_{o}}(x-x_{o})$ (1)
-where $f$ is the linear interpolation function, $x$ the target value and the subscripts 00 and $1$ represent respectively the previous and next value from target value.
-This approach ensured that the imputed values preserved the continuity and trends in the time series without introducing noticeable distortions or flattening in the data sequences.
-The rolling mean technique was applied in cases where data exhibited more noise or short-term fluctuations, allowing us to smooth these variations while still accounting for missing values.
-2.1.2 Normalization
-All features were normalized using the min-max method.
-This scaling brings all values into the range $[0,1]$, bringing all the variables expressed in different units to the same scale and making them comparable (Alaimo & Seri, 2021).
-2.1.3 Train-Test Split
-To prepare the data for modeling, each seasonal subset of the dataset was divided into training and testing sets while preserving the chronological order.
-This approach ensured that the integrity of the time series analysis was maintained, as no future information was inadvertently introduced into the training process.
-The first 80% of observations within each season were allocated to the training set, while the remaining 20% were used for testing.
-The resulting date ranges for the training and testing sets in each seasonal period are as follows:
-  • 1.
-    Winter:
-    • (a)
-      Training: from 2020-01-09 16:30:00 to 2020-02-19 04:00:00
-    • (b)
-      Testing: from 2020-02-19 04:15:00 to 2020-02-29 07:00:00
-  • 2.
-    Summer:
-    • (a)
-      Training: from 2020-07-01 14:30:00 to 2020-08-19 02:45:00
-    • (b)
-      Testing: from 2020-08-19 03:00:00 to 2020-08-31 06:00:00
-  • 3.
-    Autumn:
-    • (a)
-      Training: from 2020-10-09 22:00:00 to 2020-11-20 13:00:00
-    • (b)
-      Testing: from 2020-11-20 13:15:00 to 2020-11-30 22:45:00
-2.2 Stationarity
-In order to make statistical inference based on time series data, it is essential to ensure that the underlying data generative process fulfill the stationary property (Kirchgässner et al., 2012).
-Stationarity means that the statistical properties of the time series, such as mean and variance, remain consistent over time.
-This consistency is crucial because it ensures that the patterns observed in the past data are applicable to the future.
-A time series is stationary if it meets two conditions:
-  1. 1.
-    The mean (average value) of the series is constant over time.
-  2. 2.
-    The covariance between two points depends only on the time interval between them, not on their specific positions in time.
-Stationarity implies that the series fluctuates around a constant level with consistent variation.
-This property is essential for making reliable inferences and predictions based on the data (Tsay, 2005).
-To verify that our data met these criteria, we applied the augmented Dickey-Fuller (ADF) test (Cheung & Lai, 1995).
-The ADF test results indicate that for all variables in all the three considered periods, the null hypothesis of non-stationarity is rejected, confirming that the series are stationary.
-The results of the ADF test are presented in Tables 4, 5 and 6 in Appendix A.2.1.
+    (G4_CO2) / (OUT_CO2): dry‐air CO2 mixing ratio (ppm).
+Together with the common variables (air temperature, relative humidity and broadband solar radiation, these sensors yield five external drivers (OUT_Temp, OUT_Rad, OUT_PAR, OUT_CO2, OUT_RH) and three internal drivers (G4_PAR, G4_CO2, G4_Temp [target]).
+Data cover a single autumn window and are split chronologically 80/20 for modelling (train: 9 Oct 20:00–20 Nov 11:45; test: 20 Nov 12:00–30 Nov 21:45).
+Figure 3 shows the inside-outside temperature trajectories for the autumn period in GH4.
+Figure 3: GH4 temperatures inside and outside the greenhouse for the autumn period
+The rich PAR–CO2 couplings motivate the denser edge set of Figure 4, where directional links capture how radiation drives photosynthesis, which in turn modulates both internal CO2 and latent-heat–induced humidity changes.
+Figure 4: GH4 feature-interaction graph
+Missing-value treatment
+GH2 contains 0–75 gaps per variable (winter: up to 68; autumn: 75). GH4 exhibits 273 missing points for every external variable and for G4_PAR / G4_CO2, and 274 for G4_Temp and G4_RH.
+Gaps are imputed with a directional five-point mean: for a missing stamp at time $t$ (e.g.
+13 Sep 01:00) we average the same quarter-hour on the two previous days ($t-24$ h, $t-48$ h) and the two following days ($t+24$ h, $t+48$ h).
+This approach ensured that the imputed values preserved the continuity in the time series without introducing noticeable distortions or flattening in the data sequences.
+Normalisation and stationarity
+All features are scaled to $[0,1]$ using min–max normalisation [24].
+Augmented Dickey–Fuller tests confirm stationarity for every variable (all $p<0.01$), so no further differencing is applied [25, 26, 27].
 3 Methodology
@@
 Deep Neural Networks (DNNs) extend NNs by incorporating multiple hidden layers, enabling the modeling of complex, nonlinear relationships in data.
-Activation functions such as ReLU, sigmoid, and tanh introduce nonlinearity, while backpropagation adjusts the weights based on the error rate obtained in previous epochs (Nielsen, 2015).
+Activation functions such as ReLU, sigmoid, and tanh introduce nonlinearity, while backpropagation adjusts the weights based on the error rate obtained in previous epochs [28].
 This section outlines the deep learning techniques employed to develop predictive models of greenhouse temperature.
@@
 3.1 Recurrent neural networks
-Recurrent Neural Networks (RNNs) are specialized neural networks designed to process sequential data, making them highly suitable for time-series analysis (Jain & Medsker, 1999).
+Recurrent Neural Networks (RNNs) are specialized neural networks designed to process sequential data, making them highly suitable for time-series analysis [29].
 Unlike traditional feedforward neural networks, RNNs have recurrent connections that allow information to persist across time steps, enabling the network to capture temporal dependencies.
@@
 The hidden state $\bm{h}_{t}$ at time step $t$ is updated based on the current input $\bm{x}_{t}$ and the previous hidden state $\bm{h}_{t-1}$:
-$\bm{h}_{t}=\tanh(\bm{W}_{h}\bm{x}_{t}+\bm{U}_{h}\bm{h}_{t-1}+\bm{b}_{h})$
+$\bm{h}_{t}=\tanh(\bm{W}_{h}\bm{x}_{t}+\bm{U}_{h}\bm{h}_{t-1}+\bm{b}_{h})$ (1)
 The output $\hat{y}_{t}$ at time step $t$ is produced from the hidden state $\bm{h}_{t}$:
-$\hat{y}_{t}=\bm{W}_{y}\bm{h}_{t}+\bm{b}_{y}$
+$\hat{y}_{t}=\bm{W}_{y}\bm{h}_{t}+\bm{b}_{y}$ (2)
 where $\bm{W}_{h}$, $\bm{U}_{h}$, and $\bm{W}_{y}$ are weight matrices.
@@
 $\tanh$ is the hyperbolic tangent activation function introducing nonlinearity.
-In deeper RNNs, multiple RNN layers are stacked, passing the hidden state from one layer to the next, allowing the model to capture more complex temporal patterns (Nketiah et al., 2023).
-RNNs are effective for modeling how current and past environmental conditions influence future states, making them ideal for predicting variables like greenhouse temperature (Rodriguez et al., 1999).
-A known challenge with RNNs is the vanishing gradient problem, where gradients diminish during backpropagation through time, making it difficult to learn long-term dependencies (Behrang et al., 2010).
-However, since our data exhibits short-term dependencies, this issue do not impact our study (Dubinin & Effenberger, 2024).
+In deeper RNNs, multiple RNN layers are stacked, passing the hidden state from one layer to the next, allowing the model to capture more complex temporal patterns [30].
+RNNs are effective for modeling how current and past environmental conditions influence future states, making them ideal for predicting variables like greenhouse temperature [31].
+A known challenge with RNNs is the vanishing gradient problem, where gradients diminish during backpropagation through time, making it difficult to learn long-term dependencies [32].
+However, since our data exhibits short-term dependencies, this issue do not impact our study [33].
 3.1.1 RNN implementation
 We begin by preparing the dataset, where each input sequence $\bm{X}^{(i)}$ consists of historical environmental measurements over a period of $T$ time steps.
-The features include greenhouse temperature and relative humidity, as well as outdoor environmental variables such as temperature, relative humidity, solar radiation, and wind speed.
 The target value $y^{(i)}$ is the greenhouse temperature at the next time step.
@@
 A Dense layer with one unit at the end of the network outputs the predicted temperature.
-The model is compiled using the Adam optimizer (Kingma & Ba, 2014), which adapts the learning rate during training for efficient convergence.
+The model is compiled using the Adam optimizer [34], which adapts the learning rate during training for efficient convergence.
 The loss function is set to Mean Squared Error (MSE), suitable for regression tasks.
@@
 To model the spatio-temporal dynamics of the greenhouse microclimate, we employ directed STGNN.
-This approach integrates Graph Attention Networks (GATs) for spatial feature extraction (Veličković et al., 2018) and Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) for temporal sequence modeling.
+This approach integrates Graph Attention Networks (GATs) for spatial feature extraction [3] and Long Short-Term Memory (LSTM) networks [35] for temporal sequence modeling.
 Unlike traditional RNN-based models, which focus primarily on temporal sequences, STGNNs leverage Graph Neural Networks (GNNs) to incorporate topological and relational information among variables.
-This is well-aligned with recent advances in GNNs for time series (Jin et al., 2024).
+This is well-aligned with recent advances in GNNs for time series [36].
 By representing each environmental variable, such as internal temperature, humidity, and external factors like solar radiation and wind speed, as a node in a directed graph, we capture how changes in one variable may propagate to others through directional edges.
-Figure 4 illustrates this architecture.
+Figure 2 and 4 illustrates this architecture.
 Each node corresponds to an environmental variable, while edges encode the direction of influence.
@@
 The GAT layers learn to weigh the importance of edges and neighboring nodes, revealing spatial dependencies, while the LSTM layers capture long-term temporal patterns.
+More specifically, we adopt the GAT because its edge-wise attention coefficients are computed for each directed edge, allowing the model to weight asymmetric causal links (e.g., OUT_PAR $\rightarrow$ G4_PAR) that undirected GCN or GraphSAGE aggregators would blur.
+For the temporal branch we choose an LSTM cell: it offers the same gating mechanism as the Simple-RNN baseline (ensuring a fair comparison), remains numerically stable over our 96-step sequences, and is a well-established benchmark for reproducibility.
+It is important to emphasise that the graph is topological in feature space rather than in physical space: each node represents an environmental variable (e.g., PAR, CO2, humidity), not a sensor location, and edges encode functional or causal dependencies among variables rather than geographical proximity.
 3.2.1 Graph Attention Networks and Directed Graphs
-The Graph Attention Network (GAT) layer (Veličković et al., 2018) is employed to handle the directed graph and assign different weights to incoming edges through an attention mechanism.
+The Graph Attention Network (GAT) layer [3] is employed to handle the directed graph and assign different weights to incoming edges through an attention mechanism.
 This mechanism allows each node to weigh the importance of its neighbors’ features when updating its own representation.
 For each node $i$ and its incoming neighbor $j$, the attention coefficient $\alpha_{ij}$ is computed to determine the importance of node $j$’s features to node $i$:
-$e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{W}\bm{x}_{i}\,\|\,\bm{W}\bm{x}_{j}\right]\right)$
-$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$
+$e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{W}\bm{x}_{i}\,\|\,\bm{W}\bm{x}_{j}\right]\right)$ (3)
+$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$ (4)
 where:
@@
 The updated feature vector $\bm{h}_{i}$ for node $i$ is computed by aggregating the transformed features of its neighbors, weighted by the attention coefficients:
-$\bm{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\bm{W}\bm{x}_{j}\right)$
+$\bm{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\bm{W}\bm{x}_{j}\right)$ (5)
 where $\sigma$ is an activation function, such as ReLU.
@@
 3.2.2 Long Short-Term Memory (LSTM) Networks
-Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) are a specialized type of RNN architecture designed to better capture and retain long-term dependencies in sequential data.
+Long Short-Term Memory (LSTM) networks [35] are a specialized type of RNN architecture designed to better capture and retain long-term dependencies in sequential data.
 Unlike traditional RNN units, which can suffer from the vanishing gradient problem, LSTM cells incorporate gating mechanisms, namely the input, forget, and output gates, allowing the network to selectively add or remove information from its internal cell state.
@@
 To avoid data leakage, the feature of the target node (greenhouse temperature) at the current time step is set to zero during input preparation.
-The model is compiled using the Adam optimizer (Kingma & Ba, 2014) and the Mean Squared Error (MSE) loss function.
+The model is compiled using the Adam optimizer [34] and the Mean Squared Error (MSE) loss function.
 It is trained over $E_{p}$ epochs using batches of size $B$.
@@
 4 Analysis and results
-This section presents the results obtained from applying the RNN and directed STGNN models to the greenhouse temperature prediction task.
-Both models were trained and tested using a sequence length of $T=96$ time steps, corresponding to a full 24-hour cycle of observations recorded at 15-minute intervals.
-This setup allowed the models to capture daily patterns and short-term dependencies effectively.
-In addition, both models were trained for 32 epochs with a batch size of 96, aligning with practices in comparable studies (Belhaj Salah & Fourati, 2021; Dae-Hyun et al., 2020).
-Furthermore, 10% of the training set in each seasonal period was used as a validation set to monitor the models’ performance during training.
-This approach prevented overfitting, evidenced by the validation loss never surpassing the training loss throughout the process.
-To facilitate direct comparisons between the RNN and STGNN models, the same evaluation metrics were employed.
-Mean Squared Error (MSE) and Root Mean Squared Error (RMSE) measured the prediction error magnitude, while the coefficient $R^{2}$ assessed the proportion of variance in greenhouse temperature explained by the model’s predictions.
-4.1 Recurrent neural networks application
-Table 2 summarizes the test set results obtained by the RNN model for the three considered seasonal periods.
-Table 2: RNN test performance across different seasons
-Season MSE RMSE $\bm{R^{2}}$
-Summer 0.0032 0.056 0.937
-Winter 0.0005 0.022 0.985
-Autumn 0.0007 0.026 0.983
-The RNN demonstrates strong predictive capabilities, achieving high $R^{2}$ values across all seasons.
-Its performance is particularly impressive in winter, where $R^{2}\approx 0.985$ indicates that the model captures nearly all the variance in greenhouse temperature.
-This result depends by several factors related to greenhouse operation during the cold season.
-First, despite winter typically exhibiting large external temperature fluctuations, the greenhouse’s effectively maintains stable internal conditions, making the system’s behavior more predictable.
-Second, the reduced solar radiation intensity in winter means the greenhouse is less subject to sudden heat gains that could perturb the internal environment.
-Third, the thermal screen’s regular nighttime deployment (activated whenever the outside temperature is 2°C lower than the heating setpoint) creates a more controlled environment with reduced thermal exchanges.
-These factors combine to create a more deterministic system behavior that the RNN can effectively learn and predict, despite the external environment’s inherent variability.
-In contrast, while the summer period still exhibits a high $R^{2}$ of about $0.937$, the prediction errors (MSE and RMSE) are relatively higher, with distinct patterns visible in the model’s performance.
-As shown in Figure 9 the model’s predictions deviate significantly from observations at higher temperatures, producing a characteristic box-shaped pattern in the scatter plot.
-This behavior directly results from the greenhouse’s cooling system activation, which triggers when internal temperature reaches a threshold of 25°C.
-Since our empirical model does not explicitly account for this control system, it fails to capture the sudden temperature regulation that occurs above this threshold.
-However, examining the lower portion of the scatter plot (below normalized temperature values of 0.6), where the cooling system remains inactive, reveals that the model maintains excellent predictive accuracy.
-This dichotomy in performance clearly demonstrates how the absence of cooling system dynamics in our model impacts its predictive capabilities, particularly during periods of peak thermal stress when temperature control mechanisms are most active..
-Autumn performance falls between these two cases, with $R^{2}\approx 0.983$, indicating strong predictive accuracy albeit slightly less than that seen in winter.
-Figure 5 provides a scatter plot comparing observed and predicted greenhouse temperatures during the winter period.
-Points closely aligned along the 1:1 line signify accurate predictions.
-Figure 6 shows the time series of actual versus predicted values, illustrating how well the RNN’s predictions track the observed temperatures over time.
-Figure 5: Scatter plot of the observed vs. predicted greenhouse temperature for the winter period using the RNN model.
-Figure 6: Time series of the observed and predicted greenhouse temperature for the winter period using the RNN model.
-Similar plots for summer and autumn are included in Appendix A.1.1.
-These additional figures reveal analogous tendencies: the model accurately tracks observed temperatures, especially in winter, where conditions appear more stable, allowing the RNN to predict internal temperature with remarkable confidence.
-While the RNN achieves robust performance in all three periods, the differences across seasons reveal important insights about greenhouse control systems and model limitations.
-The near-ideal predictions in winter reflect the greenhouse’s effective temperature stabilization, creating a more predictable environment despite external variability.
-Autumn shows similar high performance, as temperatures rarely reach extremes requiring intensive control interventions.
-However, summer predictions show a clear dichotomy: excellent accuracy below the cooling threshold temperature, but significant deviations when the cooling system activates at 25°C.
-This pattern demonstrates that our empirical model’s performance is strongly tied to the greenhouse’s control regime rather than external conditions alone, performing best when passive thermal regulation dominates and struggling when active cooling interventions occur.
-4.2 Spatio-Temporal Graph Neural Network application
-Table 3 presents the directed STGNN test set results for the three seasonal periods, using the same evaluation metrics as applied to the RNN model.
-As before, the sequence length, training epochs, batch size, and validation strategy remain consistent.
-Table 3: STGNN test performance across different seasons
-Season MSE RMSE $\bm{R^{2}}$
-Summer 0.0061 0.078 0.887
-Winter 0.0023 0.048 0.947
-Autumn 0.0034 0.058 0.904
-While the STGNN model still achieves relatively high $R^{2}$ values, its overall performance is lower than that of the RNN.
-In winter, the STGNN attains an $R^{2}$ of approximately 0.947, which, although strong, falls short of the near-perfect predictions obtained by the RNN.
-Similarly, in summer and autumn, the STGNN yields lower $R^{2}$ scores and higher MSE and RMSE values than the RNN, indicating that incorporating spatial dependencies and directionality did not enhance predictive accuracy in this particular setting.
-Figure 7 provides a scatter plot of observed versus predicted temperatures for the winter period, and Figure 8 shows the corresponding time series comparison.
-These visualizations confirm that, while the STGNN’s predictions track the overall trends, the model does not reach the same level of precision demonstrated by the RNN.
-Figure 7: Scatter plot of observed vs. predicted greenhouse temperature for the winter period using the STGNN model.
-Figure 8: Time series of observed and predicted greenhouse temperature for the winter period using the STGNN model.
-Plots for summer and autumn are included in Appendix A.1.2.
-They similarly show that the STGNN model, despite considering spatial relationships and directed interactions, does not outperform the simpler temporal approach employed by the RNN.
-This outcome suggests that, for the given data and conditions, capturing complex spatial-temporal dependencies does not necessarily translate into improved predictive accuracy.
+Both architectures were trained and tested with a sequence length of $T=96$ time steps (a full 24-h cycle at 15-min resolution).
+They share identical hyper-parameters: 32 epochs, batch size 96, Adam optimiser, 10 % validation split, to ensure that any performance difference stems from the modelling paradigm rather than tuning effort [37, 38].
+For the STGNN we use a four-head GAT layer ($K=4$), concatenating the heads so that each node outputs $H\times K$ features before the LSTM.
+To compare the simple six-variable greenhouse (GH2) with the richer eight-variable PV house (GH4) we centre the discussion on the Autumn periods.
+Summer and Winter results for GH2 are presented in Appendix A.
+GH2 (2020) versus GH4 (2024)
+Table 1: Test performance in Autumn ($T\!=\!96$, 32 epochs, batch 96)
+Greenhouse Model MSE RMSE $\bm{R^{2}}$
+GH2 (2020) RNN 0.0015 0.0393 0.958
+STGNN 0.0042 0.0650 0.884
+GH4 (2024) RNN 0.0110 0.1048 0.740
+STGNN 0.0041 0.0641 0.903
+  • 1.
+    GH2 (2020): The vanilla RNN attains near-perfect accuracy ($R^{2}=0.958$) on the relatively simple six-variable setting, comfortably outperforming the directed STGNN.
+Here the microclimate is governed mainly by smooth external forcing; explicit directional edges add little beyond the temporal cues already captured by the RNN.
+  • 2.
+    GH4 (2024): Adding PAR and CO2 sensors increases both dimensionality and interaction complexity; the STGNN now surpasses the RNN by about 16 percentage points in $R^{2}$ (0.905 vs 0.740) while halving the MSE.
+Attention heads exploit directed links such as OUT_PAR $\rightarrow$ G4_PAR and G4_PAR $\rightarrow$ G4_Temp, which a purely temporal model cannot encode explicitly.
+GH2’s variables form a nearly linear control chain (outside $\rightarrow$ inside) with limited cross-talk.
+In contrast, GH4 exhibits multiple feedback loops: light drives photosynthesis, lowering internal CO2, which in turn affects transpiration and humidity, all interacting with ventilation events.
+The graph attention mechanism can re-weight these concurrent pathways on the fly, whereas the RNN must learn an entangled representation from sequences alone, hence its lower accuracy.
+For completeness, Table 2 in Appendix A reports GH2 Summer and Winter scores.
+The RNN retains a narrow edge in both cases, confirming that when the variable set is modest and control actions are well understood, a lightweight temporal network suffices.
+Figures 6(b) and8(b) visualise the GH4 test results, while Figures 5 and 7(b) provide the corresponding GH2 Autumn comparison.
+On the time series plots, inverse scaling have been performed to have Celsius degrees on the y-axis.
+Figure 5: GH2 Autumn. Observed vs. predicted temperature.
+(a) RNN scatter
+(b) STGNN scatter
+Figure 6: GH4 Autumn. Observed vs. predicted temperature.
+(a) RNN
+(b) STGNN
+Figure 7: GH2 Autumn. Time series of observed and predicted temperature.
+(a) RNN
+(b) STGNN
+Figure 8: GH4 Autumn. Time series of observed and predicted temperature.
+(a) RNN
+(b) STGNN
+Figure 6(b) contrasts the two 1:1 scatter plots for GH4.
+The STGNN cloud is tightly clustered along the diagonal, with only modest spread at the lowest temperatures, whereas the RNN points disperse into a pronounced “fan”: the network tends to over-predict cool nights and under-predict the hottest daytime spikes.
+The stacked traces in Fig.
+8(b) confirm this pattern, STGNN rises and falls almost in phase with measurements, capturing sharp midday peaks and steep nocturnal cooldowns, whereas the RNN lags and drifts, explaining its higher RMSE and lower $R^{2}$.
+In contrast, GH2 Autumn plots (Figures 5 and 7(b)) show both models hugging the diagonal and tracking the time series closely; the RNN’s slight advantage is visible only in a tighter scatter and marginally smaller peak-to-peak residuals.
+The absence of PAR/CO2 feedback loops leaves little room for the graph attention mechanism to improve upon the temporal baseline.
+Overall, these results substantiate the central claim: directional graphs pay off once sensor richness amplifies interaction pathways, while a simpler RNN already saturates performance in low-dimensional, weakly coupled settings.
 5 Discussion
-The comparative evaluation of RNN and directed STGNN models for greenhouse temperature prediction provides insights into how each model leverages available information and handles environmental complexity.
-The strong performance of the RNN, particularly during the winter period suggests that temporal patterns alone can effectively capture many of the dependencies underlying the greenhouse microclimate.
-Under these circumstances, long-term stationarity and moderate seasonal variations allow historical sequences of internal and external factors to sufficiently inform future conditions.
-As a result, the RNN consistently achieves near-perfect predictions in such stable regimes.
-However, as conditions grow more complex, the advantages of a temporal-only model appear less pronounced.
-In summer, for example, the accuracy of the RNN decreases, indicating the challenges posed by more dynamic and rapidly fluctuating environmental factors.
-Abrupt changes in solar radiation, the intermittent operation of cooling systems, or sudden shifts in external temperature highlight a broader limitation: the RNN’s inability to explicitly model the spatial and directional interactions among variables.
-While it excels in extracting temporal dependencies, it cannot directly encode the influence structure that might clarify why certain variables cause distinct, localized changes or how perturbations propagate from one factor to another.
-In principle, the STGNN architecture is designed to address this very limitation.
-By representing each variable as a node in a directed graph and capturing the directionality of interactions, the STGNN should be well-suited to handle scenarios where spatial relationships and directional effects matter.
-The GAT component of the STGNN aims to highlight influential nodes and edges, while the LSTM component extracts temporal dependencies.
-Together, this framework has the potential to disentangle complex spatio-temporal patterns that a purely temporal model may overlook.
-Yet, in our experiments, the STGNN did not outperform the RNN.
-This outcome can be attributed to several factors.
-First, the set of input variables, though representative of basic external and internal greenhouse conditions, may not be rich enough to fully leverage the STGNN’s spatial modeling capabilities.
-For instance, without additional features that intensify spatial heterogeneity (e.g., PV generation data influencing localized temperature gradients, or crop growth indicators that reflect biological processes tied to specific sections of the greenhouse), the STGNN may not find compelling reasons to depart from solutions similar to those of a temporal-only model.
-In other words, if the underlying system appears nearly homogeneous from a spatial perspective, or if directional edges do not provide sufficiently distinct pathways of influence, the added complexity of the STGNN architecture may not yield tangible benefits.
-Second, the STGNN’s performance depends on accurately learning graph structures and attention patterns that differentiate the importance of certain nodes and edges.
-If the graph connections are not strongly differentiated, the GAT may not produce markedly different node embeddings, diminishing the model’s advantage.
-The LSTM layer, while capable of capturing temporal information, cannot compensate if the spatial relationships remain underexploited.
-These observations guide the interpretation of our results.
-The theoretical superiority of STGNN in handling complex and directed spatio-temporal dependencies is not dismissed by the current findings.
-Instead, the experiments suggest that to unlock the STGNN’s full potential, the modeling scenario must present more pronounced spatial variations, clearer directional influences, or a richer set of variables that intensify these spatial-temporal complexities.
-Under such enriched conditions, the STGNN may provide insights and accuracies that temporal-only models cannot match.
+The autumn experiments on two greenhouses of increasing dimensionality illuminate when a purely temporal network is sufficient and when an explicitly directed graph becomes more useful.
+In the six-variable GH2 data set, the vanilla RNN already explains more than 95% of the temperature variance and leaves little residual structure for a graph network to exploit.
+The microclimate can be approximated as a nearly one-way control chain in which external temperature and radiation drive internal conditions with limited feedback, so the chronological context contained in 96 time steps is enough for the RNN to learn the underlying dynamics.
+By contrast, the eight-variable PV house (GH4) adds PAR and CO2 sensors that introduce two new, strongly coupled, feedback loops: (i) light raises PAR inside the house, boosting photosynthesis and thus lowering internal CO2; (ii) ventilation events triggered by heat or humidity instantaneously mix outside and inside air, altering both CO2 and RH.
+These loops generate temperature excursions that depend not only on the magnitude of each driver but on their direction and relative timing.
+The GAT layer can re-weight the edges OUT_PAR,$\rightarrow$,G4_PAR, G4_PAR,$\rightarrow$,G4_Temp and G4_CO2$\leftrightarrow$,G4_Temp on the fly, whereas the RNN is forced to embed all interactions into a single recurrent state.
+The result is a significant gain in $R^{2}$ and drop in MSE for the STGNN, confirming that directional graphs pay off once sensor richness amplifies interaction pathways.
+Several practical lessons follow.
+First, model choice should be driven by anticipated variable coupling rather than by the generic appeal of “spatial” deep learning: if only a handful of weakly interacting drivers are measured, a lightweight RNN is likely to saturate performance.
+Second, the incremental cost of a graph network is modest, one additional GAT layer with $K=4$ heads adds fewer than 30 k parameters, so upgrading to an STGNN is advisable whenever new sensors or control rules create non-trivial feedback.
+Third, graph quality matters: the manually defined edges used here relied on fundamental thermodynamic reasoning and produced stable training; future deployments should test data-driven edge discovery or adaptive pruning to avoid spurious links when variable sets grow further.
+Limitations remain.
+The directed graphs capture variable-to-variable causality but not spatial heterogeneity within the greenhouse (e.g.
+temperature gradients along the crop rows) nor operational events such as window openings or heating pulses, which would require either additional sensors or hybrid physical–empirical coupling.
+Moreover, even the richer GH4 data span only six weeks; longer records across seasons are needed to quantify generalisation.
 6 Conclusions and future developments
-This study examined the performance of Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs) in predicting greenhouse temperature under different seasonal conditions.
-Despite the STGNN’s conceptual ability to model spatial and directional relationships, the RNN model yielded higher predictive accuracy given the current set of input variables and operating conditions.
-The strong performance of the RNN, particularly during the winter period, illustrates that purely temporal modeling can effectively capture short-term dependencies.
-The implications of this research extend far beyond technical model comparisons, touching on critical aspects of sustainable agriculture and efficient land use.
-Predicting greenhouse microclimate accurately is a crucial first step toward integrating photovoltaic (PV) systems into greenhouses.
-This integration promises to optimize land use by combining energy production with food cultivation, thereby enhancing sustainable agricultural practices.
-It is instrumental in promoting sustainable land management practices that are vital in addressing global environmental challenges, such as resource efficiency and climate change mitigation.
-Looking ahead, a more diverse set of input variables may enhance the potential benefits of incorporating spatial and directional information.
-Despite the STGNN’s current performance not surpassing the RNN in this specific scenario, its architectural advantages position it as a promising framework for future greenhouse modeling challenges.
-As we move toward integrating photovoltaic systems, the STGNN’s ability to handle directed relationships will become crucial for modeling how PV panels affect local shading patterns, temperature gradients, and energy flows.
-The graph structure naturally accommodates the addition of new nodes representing PV generation parameters, inverter efficiency, and panel temperature, along with their complex interactions with the greenhouse environment.
-Similarly, when incorporating crop models, the STGNN can represent different plant varieties as separate nodes, each with unique growth parameters and responses to environmental conditions, enabling simulation of mixed-crop scenarios within the same greenhouse.
-This extensibility last to potential future additions such as energy storage systems, artificial lighting, or multiple greenhouse compartments with distinct microclimates.
-Furthermore, the attention mechanism in GAT layers could prove valuable in identifying which relationships become more or less important under different operating conditions or seasonal changes, providing insights that could inform greenhouse design and control strategies.
-This scalability and interpretability make the STGNN framework particularly valuable for developing comprehensive digital twins of increasingly complex agricultural systems, even if simpler models may suffice for basic temperature prediction tasks.
-Such tests could confirm whether additional complexity and diversity in data naturally lead to scenarios where spatial-temporal modeling offers clear advantages, ultimately guiding the development of more robust and universally applicable predictive solutions for controlled-environment agriculture.
-This study represents an initial step in the broader REGACE project’s vision of developing comprehensive digital twins for PV-integrated greenhouses.
-While our current models focus on temperature prediction, they lay the groundwork for REGACE’s more ambitious goals.
-The project aims to create an integrated modeling framework that combines greenhouse microclimate prediction with PV system performance and crop growth dynamics.
-Future work within REGACE will focus on expanding these models to incorporate PV panel characteristics, testing different panel configurations and their impact on both energy generation and crop growth.
-Additionally, the project will explore how these empirical models can complement detailed physical simulations, potentially offering rapid preliminary assessments for greenhouse design optimization.
-By developing these tools within the REGACE framework, we aim to support the widespread adoption of agrivoltaic greenhouse systems while ensuring optimal conditions for both energy generation and crop production.
-7 Acknowledgment
-This research was carried out within the framework of the European REGACE project, funded by the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101096056.
-Appendix A Supplementary Information
-A.1 Figures
-A.1.1 RNN results figures
-The graphs of the RNN results for the summer and autumn periods follow.
-Figure 9: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the RNN model.
-Figure 10: Time series of the observed and predicted greenhouse temperature for the summer period using the RNN model.
-Figure 11: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the RNN model.
-Figure 12: Time series of the observed and predicted greenhouse temperature for the autumn period using the RNN model.
-A.1.2 STGNN temperature figures
-The graphs of the directed STGNN results for the summer and autumn periods follow.
-Figure 13: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the STGNN model.
-Figure 14: Time series of the observed and predicted greenhouse temperature for the summer period using the STGNN model.
-Figure 15: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the STGNN model.
-Figure 16: Time series of the observed and predicted greenhouse temperature for the autumn period using the STGNN model.
-A.2 Supplementary tables
-A.2.1 ADF test results tables
-Hypotheses
-  • 1.
-    $\mbox{H}_{0}$: The series has a unit root (non-stationary),
-  • 2.
-    $\mbox{H}_{1}$: The series is stationary.
-Table 4: ADF test results for stationarity for the summer period
-Variable Dickey-Fuller Statistic p-value Conclusion
-G2_temp -15.855 <0.01 Stationary
-G2_RH -11.471 <0.01 Stationary
-OUT_temp -21.311 <0.01 Stationary
-OUT_RH -16.621 <0.01 Stationary
-OUT_rad -21.836 <0.01 Stationary
-OUT_wind_speed -15.131 <0.01 Stationary
-Table 5: ADF test results for stationarity for the autumn period
-Variable Dickey-Fuller Statistic p-value Conclusion
-G2_temp -16.11 <0.01 Stationary
-G2_RH -13.328 <0.01 Stationary
-OUT_temp -14.324 <0.01 Stationary
-OUT_RH -15.253 <0.01 Stationary
-OUT_rad -17.418 <0.01 Stationary
-OUT_wind_speed -10.395 <0.01 Stationary
-Table 6: ADF test results for stationarity for the winter period
-Variable Dickey-Fuller Statistic p-value Conclusion
-G2_temp -18.623 <0.01 Stationary
-G2_RH -14.903 <0.01 Stationary
-OUT_temp -13.869 <0.01 Stationary
-OUT_RH -13.673 <0.01 Stationary
-OUT_rad -17.356 <0.01 Stationary
-OUT_wind_speed -9.815 <0.01 Stationary
+This study compared Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs) for quarter-hour greenhouse, temperature prediction.
+Two data sets of growing dimensionality were analysed: a six-variable 2020 installation (GH2) and an eight-variable 2024 PV house (GH4).
+When drivers form an almost one-way chain, as in GH2, a vanilla RNN explains more than 95% of the variance and the extra graph structure yields no gain.
+In GH4, however, the addition of PAR and CO2 sensors creates coupled feedback loops that a purely temporal model cannot disentangle.
+Introducing one GAT layer with $K=4$ heads raises the STGNN to $R^{2}=0.905$ and cuts the MSE by half, while adding only approximately 30k parameters, negligible overhead by modern deep-learning standards.
+Thus, graph-based models become worthwhile once richer sensing introduces directional cross-talk, whereas a plain RNN suffices for low-dimensional, weakly coupled settings.
+The practical takeaway is clear: model choice should reflect expected coupling complexity, not architectural fashion.
+Installations with a few loosely interacting variables need nothing more elaborate than an RNN.
+Building on the insights gained from a single geographic region, future work will actively extend this analysis across diverse climates, greenhouse designs, and sensor setups to robustly demonstrate the broader applicability and scalability of the STGNN approach.
+This work is an early step toward the REGACE project’s digital-twin vision for PV-integrated greenhouses.
+Although we focus here on temperature, the graph framework readily accommodates new nodes.
+Future work will embed PV-panel characteristics, alternative panel layouts, and crop-growth variables, assessing their joint impact on energy yield and microclimate.
+Coupling these empirical models with detailed physical simulations will enable rapid design screening and, ultimately, integrated control of agrivoltaic greenhouses, supporting wider adoption while ensuring optimal conditions for both power generation and crop production.
+CRediT authorship contribution statement
+Emiliano Seri: Writing – original draft & editing, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization.
+Marcello Petitta: Writing – original draft & editing, Conceptualization, Investigation, Formal analysis, Methodology, Resources, Funding acquisition, Supervision.
+Chryssoula Papaioannou: Data curation.
+Nikolaos Katsoulas: Data curation.
+Cristina Cornaro: Writing – original draft & editing, Conceptualization, Formal analysis, Funding acquisition, Supervision.
+Data availability
+The authors do not have permission to share data.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgments
+This research was carried out in the framework of the REGACE project funded by the European Union under Grant Agreement No 101096056.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or CINEA.
+Neither the European Union nor the granting authority can be held responsible for them.
+Appendix A GH2 Summer and Winter results
+A.1 Test metrics
+Table 2: GH2 Summer and Winter. Test-set performance ($T\!=\!96$, 32 epochs, batch 96)
+Season Model MSE RMSE $\bm{R^{2}}$
+Winter RNN 0.0019 0.0433 0.957
+STGNN 0.0019 0.0441 0.955
+Summer RNN 0.0061 0.0781 0.883
+STGNN 0.0062 0.0785 0.882
+A.2 Winter figures
+Figure 9: GH2 temperatures inside and outside the greenhouse for the winter period
+Figure 10: GH2 Winter. Observed vs. predicted temperature.
+(a) RNN scatter
+(b) STGNN scatter
+A.3 Summer figures
+Figure 11: GH2 Temperatures inside and outside the greenhouse for the summer period
+Figure 12: GH2 Summer. Observed vs. predicted temperature.
+(a) RNN scatter
+(b) STGNN scatter